OpenStudy (anonymous):

If f(x) is the height, in cm, of a sunflower plant which is x days old, which of the following statements best describes the meaning of f(60) = 210? The height of the sunflower plant is 60 cm when it is 210 days old. The height of the sunflower plant is 210 cm when it is 60 days old. The height of the sunflower plant is 210 cm when it is 3.5 days old. The height of the sunflower plant is 60 cm when it is 3.5 days old. @whatdoesthismean

OpenStudy (anonymous):

I already tried working this out before but I was stuck, is it B?

OpenStudy (anonymous):

okay since f(x) is the hieght, and it equals 210, 210 is the number of cm

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so then you can guess, that 60 is the number of days which is like you said, it is B

OpenStudy (anonymous):

YEAH! Thanks so much, I'm a 7th grader doing 9th grade work so thats why I need so much help haha

OpenStudy (anonymous):

The first five terms of a sequence are shown below. 4, 7, 10, 13, 16 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 3n - 1; n ≥ 1 f(n) = 3n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (anonymous):

This is more of trial and error.

OpenStudy (anonymous):

The first one isn't correct

OpenStudy (anonymous):

Right? :/

OpenStudy (anonymous):

SO basically you take the nth term (for example 4 is the first term, and etc.) and put it in for x. so ill do it for B. f(1). it results in 4. This works so you try it for the next nth term, which is 2.

OpenStudy (anonymous):

and yes A is 100% not correct

OpenStudy (anonymous):

And.... B isn't correct also is it?

OpenStudy (anonymous):

and since when you plug in 1 for n, C and D also do not work, those are ruled out, and you have B for your answer :)

OpenStudy (anonymous):

So yes you are correct! if you have any more questions please tag me in another post, its getting very long again lol

OpenStudy (anonymous):

Oh ok thanks, I have a few more haha you are really awesome